### v2

# MMLA: Multi-Environment, Multi-Species, Low-Altitude Drone Dataset

Abstract
Real-time wildlife detection in drone imagery supports critical ecological and conservation monitoring. However, standard detection models like YOLO often fail to generalize across locations and struggle with rare species, limiting their use in automated drone deployments. We present MMLA, a novel multi-environment, multi-species, low-altitude drone dataset collected across three sites (Ol Pejeta Conservancy and Mpala Research Centre in Kenya, and The Wilds in Ohio), featuring six species (zebras, giraffes, onagers, and African wild dogs). The dataset contains 811K annotations from 37 high-resolution videos. Baseline YOLO models show performance disparities across locations while fine-tuning YOLOv11m on MMLA improves mAP50 to 82%, a 52-point gain over baseline. Our results underscore the need for diverse training data to enable robust animal detection in autonomous drone systems.
Figure 1: Examples of the MMLA dataset and example detection results from a YOLOv11m model fine-tuned on MMLA. The dataset spans multiple species and environments, enabling robust model training for low-altitude wildlife detection.
1 Introduction
Animal ecology studies increasingly rely on vision-based technologies for efficient, noninvasive wildlife monitoring. Camera traps, high-altitude and low-altitude drones each offer distinct advantages for different research goals. Machine learning models tailored for animal ecology data can rapidly analyze these datasets to produce valuable insights [31], but their effectiveness varies significantly across monitoring approaches. While camera traps and high-altitude drone surveys have established methodologies and datasets [33, 34], low-altitude drone studies represent an untapped opportunity. These deployments offer more detailed imagery than high-altitude approaches, covering broader areas than stationary camera traps and providing valuable insights into animal interactions and decision-making [21]. Critically, low-altitude drones are increasingly deployed as autonomous systems that can actively track wildlife [18, 20], requiring models with high precision and localization.
While effective on benchmark datasets, standard computer vision models like YOLO [14] perform poorly when applied to wildlife detection in low-altitude drone imagery across varied natural environments. This performance gap stems from the specific requirements of autonomous wildlife-tracking drones: high precision detection with minimal false positives and accurate object localization (measured by IoU) are essential, as false detections or imprecise bounding boxes can cause autonomous systems to make incorrect flight decisions [20]. Additionally, environmental variability across field sites introduces substantial domain shifts, causing models trained in one location to fail when deployed elsewhere.
To address this gap, we present a novel multi-environment, multi-species, low-altitude drone dataset (MMLA), the first low-altitude dataset explicitly designed for developing robust computer vision models for autonomous wildlife-tracking drones. Our dataset spans three distinct locations (Ol Pejeta Conservancy and Mpala Research Centre in Kenya and The Wilds Conservation Center in Ohio, USA) and six species: plains zebras, Grevy’s zebras, reticulated and Masai giraffes, Persian onagers, and African wild dogs. MMLA contains 155K frames (1 hr, 26 min), with manually annotated bounding boxes around each animal optimized for training high-precision detection models.
The creation of MMLA represents a significant investment in advancing wildlife conservation technology, involving over 150 person-hours of fieldwork across two continents. By releasing this dataset and fine-tuned model weights, we invite the computer vision community to tackle these challenging real-world ecological problems and advance the state of the art in autonomous wildlife monitoring technology.
2 Background & Related Work
Feature Camera High-alt. Low-alt.
Trap Drone Drone
Size1 med-large v. small med
(128-256) (0-16) (64-128)
Processing Post-hoc Post-hoc Real-time
batches batches video
Metric Low false F1-score High prec.
negative & mAP50
Vision task Detect, Counting, Localization
Species ID Species ID
Models MegaDetector POLO [24], None
[3] HerdNet [9]
Table 1: Comparison of vision-based animal ecology monitoring approaches. Low-altitude drone studies differ from camera trap and high-altitude drone studies for processing, relevant metrics, and computer vision tasks. There are no pre-trained models aimed at low-altitude drone studies.
1Typical animal size in pixels [12].
2.1 Monitoring Approaches for Animal Ecology
Figure 2: Animal imagery collected via (a) camera traps [33], (b) high-altitude drone [13], and (c) low-altitude drone missions [15].
Vision-based wildlife monitoring approaches differ significantly in their technical specifications, applications, and computer vision requirements (Tab. 1 and Fig. 2). Camera traps capture medium to large-sized objects (128-256 pixels) [12] and excel at species identification but are stationary. High-altitude drones capture very small objects (0-16 pixels) [34] and are optimized for population counts across large areas but lack detail for individual ID [29] or behavioral analysis [15]. Low-altitude drone studies capture medium-sized objects (100 pixels) [15], offering a balance that enables behavioral analysis, individual ID, and tracking studies. The optimal resolution for these ecological studies is approximately 100-300 pixels per animal [20], which low-altitude drones provide consistently. Unlike camera traps and high-altitude drones, which typically process data in post-hoc batches, low-altitude drone applications often require real-time processing to support autonomous flight decisions [18, 20]. These fundamental differences in monitoring approaches have driven the development of specialized computer vision models for camera traps (e.g., MegaDetector [3]) and high-altitude imagery (e.g., HerdNet [9], POLO [24]). However, no pre-trained models specifically address the unique challenges of low-altitude drone imagery.
2.2 Existing Wildlife Drone Datasets
Dataset Categories Image Count Size1 Multi-Location2
Population estimates [10] elephant, zebra, giraffe 561 RGB 50
Bird detection [35] bird 24K RGB 35
Seabird colonies [11] albatross, penguin 4K RGB 300
Conservation Drones [5] human, elephant, lion, dog 48 thermal videos 35
Group behavior [21] zebra, gazelle, waterbuck, buffalo, gelada, human 2K RGB 56
KABR [15] giraffe, zebra 130K videos 100
WAID [25] sheep, cattle, seal, camel, kiang, zebra 14K RGB 166
WildLive [8] zebra, giraffe, elephant 19K RGB 100
WildlifeMapper [22] zebra, hartebeest, cattle, shoats, zebra 11K RGB
BuckTales [27] antelope 22.5K RGB 1000
MMLA (ours) zebra, onager, giraffe, African wild dog 155K RGB 100
1Size indicates the typical animal size in pixels. 2Multi-Location indicates datasets capturing the same species across multiple geographic locations.
Table 2: Summary of multi-species aerial datasets. Note that MMLA has large animals (100px) and multiple locations, unlike prior work.
A critical gap exists in wildlife monitoring datasets: no multi-species, multi-location aerial footage dataset provides sufficient resolution with the same species captured across multiple environments (Tab. 2). While the global bird detection dataset [35] contains multi-location imagery across 13 ecosystems, the animals are very small (35 pixels) and limited to avian species. Existing mammal datasets [10, 15, 25, 21, 8, 22] contain thousands of annotated samples but are geographically limited to single locations for each species, creating detection models that may not transfer well across environments [6, 32]. Specialized datasets like BuckTales [27] and WildLive [8] focus on single locations or specific computer vision tasks. BuckTales provides 22.5K RGB frames of antelopes for multi-object tracking and re-identification. The WildLive dataset includes multiple species (zebras, giraffes, and elephants) but all were captured at a single location (Ol Pejeta Conservancy) [8]. These datasets highlight the importance of high-quality, ecologically relevant datasets tailored to specific computer vision problems but do not address the critical need for cross-environment generalization.
2.3 Animal Detection Models
Existing wildlife detection approaches are tailored for camera traps and high-altitude imagery, making them ill-suited for the demands of low-altitude drone studies. Camera trap pipelines like MegaDetector [3] focus primarily on reducing false negatives to eliminate blank images [33] before specialized species classification models process them. Models for high-altitude imagery use point-based approaches optimized for accurate counting from minimal pixel information [24, 9]. These methods fail to meet the high precision and accurate localization requirements of automated low-altitude drone studies. Low-altitude drones often need real-time processing with high precision metrics to guide autonomous flight decisions [18], where false positives can significantly impair performance by causing drones to track non-animal objects. Our MMLA dataset aims to bridge this gap by enabling the development of robust models that can generalize across environments and species with autonomous drones.
3 MMLA Dataset
The MMLA dataset [^1] https://huggingface.co/collections/imageomics/mmla-project contains annotated drone collected for wildlife monitoring and conservation research. The data was collected at three locations: Ol Pejeta Conservancy and Mpala Research Centre, both in Laikipia, Kenya, and The Wilds Conservation Center, in Ohio, USA. The dataset contains six species: Plains zebra (Equus quagga), Grevy’s zebra (Equus grevyi), Persian onager (Equus hemionus onager), Reticulated giraffe (Giraffa reticulata), Masai giraffe (Giraffa camelopardalis tippelskirchi), and African Wild Dog (Lycaon pictus). The variety of species and environments in the MMLA dataset enables the evaluation of model generalizability in detecting animals across different settings. The dataset includes 155,074 video frames, captured at 30 frames per second, and contains a variety of habitats (details in App. 9 Tab. 4). Each video frame has COCO format annotations (bounding boxes with normalized coordinates). The dataset contains 811K total annotations, which were completed manually by a team of experts using CVAT [7] and kabr-tools [19].
The Ol Pejeta dataset contains 29,268 frames of Plains zebras at Ol Pejeta Conservancy in Kenya (Jan-Feb 2025), collected using the autonomous WildWing system [18] across seven video sessions during the 2025 WildDrone Hackathon [36]. The Mpala dataset contains 104,062 frames of reticulated giraffes, Grevy’s and Plains zebras at Mpala Research Center in Kenya (Jan 2023), collected manually using a DJI Air 2S drone across five sessions [15]. The Wilds dataset contains 21,744 frames of African wild dogs, Persian onagers, reticulated and Masai giraffes and Grevy’s zebras at The Wilds Conservation Center in Ohio (Apr-Jul 2024), collected using both DJI Mini and Parrot Anafi drones, with with a combination of autonomous WildWing system [16] and manual collection.
Model IoU (%) Precision (%) Recall (%) F1 (%) mAP50 (%)
YOLOv5m 20.88 34.04 24.94 27.09 28.79
YOLOv8m 20.23 32.20 24.78 26.61 28.00
YOLOv11m 21.77 35.34 25.83 28.14 29.85
YOLOv11m Finetuned on MMLA 74.39 83.43 81.39 81.76 82.40
Table 3: Performance of YOLO variants on the MMLA dataset. Fine-tuning YOLOv11m yields the highest performance across all metrics.
4 Evaluation
We evaluate the performance of three popular YOLO model (YOLOv5m, YOLOv8m, and YOLOv11m) for animal detection in the MMLA dataset. The YOLO (You Only Look Once) [14] family of models represents a popular choice for animal ecology field studies due to their speed, ease of use, and accuracy [17]. We selected medium-sized variants to balance detection performance with computational efficiency, as field studies often utilize edge devices, such as laptops and Raspberry Pis, with limited memory and compute[38, 4]. YOLO models perform detection in a single forward pass, making them suitable for real-time applications on resource-constrained devices, such as those used for autonomous drone deployments [18, 8].
4.1 Pretrained YOLO
Our initial evaluations were conducted using the pre-trained models without additional fine-tuning to assess their out-of-the-box performance on our dataset. We extracted approximately 500 frames from each of the ten sampling sessions to ensure even coverage across the MMLA dataset for all the locations and species. Note: we combined the two sessions collected at the Ol Pejeta Conservancy into one session for the testing, as all videos captured Plains zebras in the same park area on consecutive days. The evaluation subset consists of 5,073 frames, of which 1,050 from Ol Pejeta Conservancy, 1,923 from Mpala Research Center, and 2,100 from The Wilds Conservation Center. See App. 10 for detailed assessment of the baseline models’ performance for each location and selected species present at most locations (zebra and giraffe).
4.2 Fine-tuned YOLOv11m
YOLOv11m achieved the best performance on the MMLA dataset, with an 18.6% F1 score and 32.8% mAP50 score, so this model was selected for fine-tuning on the dataset. We implemented a stratified data partitioning strategy to ensure robust model generalization across our multi-location wildlife imagery dataset. We divided the data into training (58%) and test (42%) splits, carefully maintaining proportional representation across all three ecological zones and five target species (reticulated and Masai giraffe, Plains zebra, Grevy’s zebra, Persian onager, and African wild dog. We assigned entire video sequences to a single partition to prevent data leakage, ensuring that temporally adjacent frames never appeared in different splits when possible. Furthermore, we regularly sampled non-consecutive frames (every 10) to reduce temporal redundancy while preserving diversity in environmental conditions, animal behaviors, and camera perspectives. This approach maintained statistical consistency across partitions while ensuring independence between the training and evaluation data, with each species and location represented in at least two partitions to facilitate a thorough assessment of the model’s cross-habitat performance. Please refer to our GitHub repository  [^2] https://github.com/Imageomics/mmla for instructions on generating the train/test split. This division of training/test data by session and location ensures that the model is evaluated on unseen environments, individuals, and contexts, making it a robust benchmark for testing generalization across ecological and geographic domains.
We trained the model for 50 epochs (approximately 2 hours) on the Ohio Supercomputer Cluster, using two Tesla V100-PCIE-16GB GPUs. The hyperparameters and optimizers were left as the default for YOLO fine-tuning using scripts provided in the ultralytics library [14]. We provide scripts to reproduce our results in our GitHub repository and provide the fine-tuned weights [^3] https://huggingface.co/imageomics/mmla. Additional training details may be found in App. 11.
5 Results
We report each model’s IoU, precision, recall, F-1, and mAP50 scores in Tab. 3. Comparing the baseline YOLOv11m performance against the fine-tuned model reveals dramatic improvements across all metrics. Fine-tuning significantly improved model performance across locations and species. At Mpala, F1 increased by 34.3% and mAP50 by 33.8%, while The Wilds showed the most dramatic gains with F1 and mAP50 rising by 90.8% and 83.5%, respectively, indicating strong generalization to this previously underperforming site. These location-specific variations highlight distributional differences between fine-tuning data and habitat features. Species-level improvements were substantial, particularly for African wild dogs (+94.9% F1), onagers (+70.2%), and giraffes (+30.4%). Even zebras, which had the highest baseline performance, improved by 22.0% F1.
Figure 3: Prediction accuracy for fine-tuned YOLOv11m across the four classes (zebra, giraffe, onager, and dog) and background. Results suggest that the model could be improved to more accurately detect the background and might be overfitted for African wild dog detection.
The normalized confusion matrix (Fig. 3) for the fine-tuned YOLOv11m model reveals strong species discrimination capabilities with minimal cross-species confusion. Zebras show minimal confusion with other classes due to their distinctive stripe patterns. Giraffes occasionally register false positives with background elements, likely due to their complex body patterns and tall stature blending with vegetation. Onagers demonstrate strong classification accuracy, while African wild dogs exhibit near-perfect classification due to their distinctive coat patterns and body shapes. Background elements occasionally trigger false positives across species, with slightly higher rates for giraffes and onagers, indicating successful learning of discriminative features while maintaining robustness for autonomous wildlife tracking applications.
6 Limitations
Our work has three main limitations that may be addressed in future work: the dataset itself, model architecture, and computational efficiency. One, our dataset has species-specific location limitations, particularly for African wild dogs, Masai giraffes, and Persian onagers, which were only present at The Wilds. Future work could combine others’ datasets (Tab. 2) with MMLA to include these species across multiple environments. Further, we combine the two zebra species and giraffe species into one class. Two, we used a baseline YOLO model without any architectural improvements specific to animal detection from drone imagery. Other works on small-object animal detection and models tailored to drone imagery (WILD-YOLO [26], WildARe-YOLO [2], and Drone-YOLO [37]), could further improve performance on the MMLA dataset. Three, our current approach does not address deployment on resource-constrained drone platforms. Future work could incorporate energy-aware algorithms that explicitly consider edge constraints, such as dynamic pruning for edge inference [28].
7 Conclusion
Low-altitude drone studies offer a valuable middle ground between the detailed imagery of camera traps and the broad coverage of high-altitude surveys, enabling new behavioral and movement studies. Autonomous drones combined with robust detection models can transform wildlife monitoring from a static, pre-planned process to an adaptive system that responds to animal movements in real-time [18]. Deploying such systems involves important tradeoffs for ecologists. Lower flight altitudes provide better resolution but may increase animal disturbance, potentially altering natural behaviors [1, 30]. The balance between human and automated analysis also presents challenges [17]. Manual analysis by trained biologists remains the reliability standard but limits study scale, while automated detection can dramatically increase processing speed but requires domain-specific training to achieve necessary reliability.
The dramatic improvements achieved through fine-tuning suggest that with appropriate dataset development, AI-assisted wildlife monitoring can achieve the reliability needed for scientific applications while significantly increasing wildlife studies’ temporal and spatial coverage. By contributing both the MMLA dataset and fine-tuned model weights to the research community, we hope to accelerate the development of robust computer vision solutions for wildlife conservation and establish a foundation for the next generation of intelligent, autonomous wildlife monitoring systems.
8 Acknowledgments
We want to thank Samuel Mutisya and William Njoroge at Ol Pejeta Conservancy for their support in facilitating the WildDrone 2025 Hackathon. This work was supported by WildDroneEU (EU Horizon Europe MSCA grant 101071224), the Imageomics Institute (NSF HDR Award 2118240), and ICICLE (NSF grant OAC-2112606).